OpenStudy (anonymous):

Create your own quadratic equation and demonstrate how it would be solved by graphing, factoring, the quadratic formula, and by completing the square.

OpenStudy (anonymous):

start of with (ax+b) * (cx+d) and expand it.

OpenStudy (anonymous):

is that the one with completing the square?

OpenStudy (anonymous):

i just need help with the graphing part, i made the equation: x^2+6x+9 and i did completing the square with y=x^2+6x-49 and i got y=1(x+3)^2+40...i'm not sure how to graph it

OpenStudy (anonymous):

it is a standard x^2 curve shifted 3 to left 40 up it seems